OpenStudy (_t_b_w_):

Help? Will medal! Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (blackfireskull):

what do you think it is? @_T_B_W_

Awolflover1 (awolflover1):

Divide both sides by -5. \[ \frac{ −5(|x+1|) }{ -5}=\frac{ 10 }{ -5}\]

OpenStudy (blackfireskull):

it's no solution.

Awolflover1 (awolflover1):

No Direct Answers

OpenStudy (_t_b_w_):

Oh sorry, I was getting something and I think that its no solutions

Awolflover1 (awolflover1):

yes you're correct @_T_B_W_

OpenStudy (_t_b_w_):

Yay! thank you :)

Awolflover1 (awolflover1):

yw